You are throwing a party, and you need 2 bags of chips for every 20 guest how many bags of chips do you need?- Equivalent ratios


1
Expert's answer
2021-04-15T06:31:15-0400

Let N=N= the number of guests, m=m= the number of bags of chips.

Then


Nm=202\dfrac{N}{m}=\dfrac{20}{2}

Solve for mm


m=N(2)20m=\dfrac{N(2)}{20}

m=N10m=\dfrac{N}{10}

We need m=N10m=\dfrac{N}{10} bags of chips for NN guests.